What happens to the confidence interval as the sample estimate increases?
Assuming that other measures remain the same, as the sample estimate increases both ends of the confidence interval will increase. In effect, the confidence interval will be translated to a higher value without any change in its size.
Assuming that other measures remain the same, as the sample estimate increases both ends of the confidence interval will increase. In effect, the confidence interval will be translated to a higher value without any change in its size.
Assuming that other measures remain the same, as the sample estimate increases both ends of the confidence interval will increase. In effect, the confidence interval will be translated to a higher value without any change in its size.
Assuming that other measures remain the same, as the sample estimate increases both ends of the confidence interval will increase. In effect, the confidence interval will be translated to a higher value without any change in its size.
How many red aces in a deck of 52 cards?
There are two. One in each suit, two red suits, hearts and diamonds.
How many questions can you miss to get an 83 on a 150 question test?
100 - 83 = 17
17/100 x 150 = 25.5 missed
with rounding you could miss 25 or 26
50%. Its the law of the land, or God.
What is the theoretical probability if you hit 20 of 32 balls at bat?
There is no information on the relevant theoretical basis for estimating the probability: only an empirical or experimental, basis.
How are going to determine the probability of an event?
See the Basic Rules for Probability section in the related link.
100 million
What is a probability that the selected card will be an ace if one is missing?
The probability of drawing a Ace from a standard deck of 52 cards if one Ace is missing is 3 in 51, or about 0.05882. If the missing card is not an Ace, then the probability is 4 in 51, or about 0.07843.
What is the Difference between fuzzy logic and probability?
The difference between probability and fuzzy logic is clear when we consider the underlying concept that each attempts to model. Probability is concerned with the undecidability in the outcome of clearly defined and randomly occurring events, while fuzzy logic is concerned with the ambiguity or undecidability inherent in the description of the event itself. Fuzziness is often expressed as ambiguity rather than imprecision or uncertainty and remains a characteristic of perception as well as concept.
Your question is a bit difficult to understand. I will rephrase it as follows:
What is the probability of getting a head if a coin is flipped once?
p = 0.5
What is the probability of getting 2 heads if a coin is flipped twice =
The possible events are HT, TH, HH, TT amd all are equally likely. So the probability of HH is 0.25.
What is the probability of getting at least on head if the coin is flipped twice.
Of the possible events listed above, HT, TH and HH would satisfy the condition of one or more heads, so the probability is 3 x 0.25 = 0.75 or 3/4.
Also, since the probability of TT is 0.25, and the probability of all events must sum to 1, then we calculate the probability of one or more heads to be 1-0.25 = 0.75
Is an open container probable cause to search?
Probable cause to search WHAT?
A vehicle? Yes, but only the passenger compartment or any area where the passengers of the vehicle might have access.
To search a building or a residence - no. Only that area which is in plain sight
What is the probability of getting heads from a coin 9 times in a row?
If you continue tossing the coin forever, it is effectively a certainty. So probability = 1
If you toss it only 9 times, it is (1/2)9 = 1/512 = 0.00195 (approx)
What is the probability of a compound event?
It depends on whether or not the events are independent.
When is probability theory used in statistics?
All the time. Statistic is based on the application of probability theory!
0.7734 is greater than 0.75 and the rest are less than 0.75
What is the probability of rolling an even number on a six sided die and then rolling a 4?
1/12, 1/2*1/6
How do i find the sum of outcomes in probability?
Select an experiment that has a random result rather than one that is deterministic. The result of the experiment is the outcome of the probabilistic experiment.
What is the probability of getting a number greater than 1 when rolling a regular 6-sided die?
The probability is 5/6 (83.333%).
its most ikely 6 to 25 or 1 to 4 or 1 to 2 or 1 or 6 to 5........i hope this helps now you have 4 choices and one could be the answer.
I disagree.
What is the probablity of winning the lottery?
Different lotteries have different schemes and so the probabilities are different.
What is the probability that the arrow will land on either Saturday or Sunday both times?
The answer depends on where the arrow is being thrown!